# On the Identifying Power of Generalized Monotonicity

### v2

@@ Appendix B Details of Examples
The identified set in (36) equals (7) for $d=0$ with $Y$ and $Z$ binary. Thus, the identified set for $\mathbb{E}_{Q}[Y_{0}]$ relative to $\mathbf{Q}$ corresponds to the identified set relative to $\mathbf{Q}^{\ast}_{3}$, the set of distributions that satisfy mean independence, 3.1. By the same sandwich argument used to prove Theorem 3.3, the identified set for $\mathbb{E}_{Q}[Y_{0}]$ relative to $\mathbf{Q}$ corresponds to the identified set relative to $\mathbf{Q}^{\ast}_{1}$, and thus imposing this ARUM has no identifying power for $\mathbb{E}_{Q}[Y_{0}]$ beyond instrument exogeneity.
Finally, we show that there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and (36) is strictly smaller than (5), so that $\Theta_{0}(P,\mathbf{Q})$ is not given by (5). We do so by providing a numerical example. Consider the $P$ specified in Table 8 and the $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ specified in Tables 9 and 10 respectively, where we write $q(y_{0}y_{1}y_{2},d_{0}d_{1})=Q\{Y_{d}=y_{d},D_{z}=d_{z},~{}(d,z)\in\mathcal{D}\times\mathcal{Z}\}$ and omit any $q(\cdot)=0$. One can check that both $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ rationalize $P$ and satisfy Assumption 2.1. One can further check that both $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ satisfy the restriction in (35), so that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Evaluating (36) at $P$ gives the identified set for $\mathbb{E}_{Q}[Y_{0}]$ relative to $\mathbf{Q}$ as $[0.2518,0.8167]$. One can further check that the two endpoints are attained by $\mathbb{E}_{Q_{\text{arum},\text{min}}}[Y_{0}]=0.2518$ and $\mathbb{E}_{Q_{\text{arum},\text{max}}}[Y_{0}]=0.8167$. On the other hand, if one evaluates (5) by setting $z^{*}(0)=0,1$ at the same $P$, the resulting bounds for $\mathbb{E}_{Q}[Y_{0}]$ equal $[p_{10|0},1-p_{00|0}]=[0.2518,0.8937]$ and $[p_{10|1},1-p_{00|1}]=[0.2372,0.8167]$ respectively. In both cases, (36) is strictly contained in (5). Hence, $\Theta_{0}(P,\mathbf{Q})$ is not given by (5).

### v3

@@ Appendix B Details of Examples
The identified set in (30) equals (8) for $d=0$ with $Y$ and $Z$ binary. Thus, the identified set for $\mathbb{E}_{Q}[Y_{0}]$ relative to $\mathbf{Q}$ corresponds to the identified set relative to $\mathbf{Q}^{\ast}_{3}$, the set of distributions that satisfy mean independence, 3.2. By the same sandwich argument used to prove Theorem 3.3, the identified set for $\mathbb{E}_{Q}[Y_{0}]$ relative to $\mathbf{Q}$ corresponds to the identified set relative to $\mathbf{Q}^{\ast}_{1}$, and thus imposing this ARUM has no identifying power for $\mathbb{E}_{Q}[Y_{0}]$ beyond instrument exogeneity.
Finally, we show that there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$ and (30) is strictly smaller than (5), so that $\Theta_{0}(P,\mathbf{Q})$ is not given by (5). We do so by providing a numerical example. Consider the $P$ specified in Table 8 and the $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ specified in Tables 9 and 10 respectively, where we write $q(y_{0}y_{1}y_{2},d_{0}d_{1})=Q\{Y_{d}=y_{d},D_{z}=d_{z},~{}(d,z)\in\mathcal{D}\times\mathcal{Z}\}$ and omit any $q(\cdot)=0$. One can check that both $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ rationalize $P$ and satisfy Assumption 2.1. One can further check that both $Q_{\text{arum},\text{min}}$ and $Q_{\text{arum},\text{max}}$ satisfy the restriction in (29), so that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Evaluating (30) at $P$ gives the identified set for $\mathbb{E}_{Q}[Y_{0}]$ relative to $\mathbf{Q}$ as $[0.2518,0.8167]$. One can further check that the two endpoints are attained by $\mathbb{E}_{Q_{\text{arum},\text{min}}}[Y_{0}]=0.2518$ and $\mathbb{E}_{Q_{\text{arum},\text{max}}}[Y_{0}]=0.8167$. On the other hand, if one evaluates (5) by setting $z^{\ast}(0)=0,1$ at the same $P$, the resulting bounds for $\mathbb{E}_{Q}[Y_{0}]$ equal $[p_{10|0},1-p_{00|0}]=[0.2518,0.8937]$ and $[p_{10|1},1-p_{00|1}]=[0.2372,0.8167]$ respectively. In both cases, (30) is strictly contained in (5). Hence, $\Theta_{0}(P,\mathbf{Q})$ is not given by (5).
@@
Appendix C Additional Examples of Models That Satisfy Assumption 2.2
In Section 4, we considered examples of restrictions on potential treatments previously considered in the literature that satisfy generalized monotonicity. We now consider three additional such examples.
Example C.1.
$D_{z}=\mathbbm{1}\{g_{10}(z)+U_{10}\geq 0\}~{}.$ (31)
Let $\overline{\mathcal{Z}}=\operatorname*{argmax}_{z\in\mathcal{Z}}\{g_{10}(z)\}$, and let $\underline{\mathcal{Z}}=\operatorname*{argmin}_{z\in\mathcal{Z}}\{g_{10}(z)\}$. Then, $Q$ satisfies Assumption 2.2 with $\mathcal{Z}^{\ast}(1)=\overline{\mathcal{Z}}$ and $\mathcal{Z}^{\ast}(0)=\underline{\mathcal{Z}}$. To contrast with Example 4.4, note that (11) holds if and only if $\overline{\mathcal{Z}}$ and $\underline{\mathcal{Z}}$ are both singletons.
Example C.2.
$Q\{D_{1}=2\mid D_{0}\neq D_{1}\}=1~{}.$ (32)
The condition in (32) states that if the choice of a family changes upon receiving a Head Start offer, then they must choose Head Start when receiving the offer. In other words, it cannot be the case that upon receiving a Head Start offer, a family switches from no preschool to preschools other than Head Start, or the other way around. Assumption 2.2 then holds with $z^{\ast}(0)=z^{\ast}(1)=0$ and $z^{\ast}(2)=1$. To see this, note that (32) implies $Q\{D_{0}\neq D_{1},D_{1}\neq 2\}=0$ and thus
$Q\{D_{0}\neq 0,D_{1}=0\}$ $=Q\{D_{0}\neq D_{1},D_{1}=0\}=0~{},$
$Q\{D_{0}\neq 1,D_{1}=1\}$ $=Q\{D_{0}\neq D_{1},D_{1}=1\}=0~{},$
$Q\{D_{1}\neq 2,D_{0}=2\}$ $\leq Q\{D_{0}\neq D_{1},D_{1}\neq 2\}=0~{}.$
Note in this example Assumption 2.2 still holds although $|\mathcal{Z}|<|\mathcal{D}|$. See Bai et al. (2025) for results on the sharp testable implications of the assumptions for this example and Example C.3.
Example C.3.
$Q\{D_{1}=1\mid D_{0}=1\}$ $=1~{},$ (33)
$Q\{D_{2}=2\mid D_{0}=2\}$ $=1~{}.$ (34)
$Q\{\mathbbm{1}\{D_{1}=2\}$ $=\mathbbm{1}\{D_{0}=2\}\mid D_{0}\neq 1,D_{1}\neq 1\}=1~{},$ (35)
$Q\{\mathbbm{1}\{D_{2}=1\}$ $=\mathbbm{1}\{D_{0}=1\}\mid D_{0}\neq 2,D_{2}\neq 2\}=1~{}.$ (36)
$Q\{D_{0}\neq 0,D_{1}=0\}=0~{}.$
$Q\{D_{0}\neq 0,D_{2}=0\}=0~{},$
$Q\{D_{1}\neq 1,D_{0}=1\}=0~{}.$
It therefore remains to argue by contradiction that
$Q\{D_{1}\neq 1,D_{2}=1\}=0~{}.$ (37)